The intent of this assignment is just basic accounting and financial statement construction These will get more complicated over the course of the semester, but this one is just to ensure that everyone is on the same page. Pay close attention to the notes at the end of the assignment and the definitions posted to the blackboard site (Ratio Definitions and Case Analysis). You will only need each number on time. Check the syllabus for the due datel Below are numbers from a balance sheet and income statement. Construct the financial ratios requested below. Sales Operating Costs Depreciation Interest Expense Tax Expense Cash Receivables Inventories Fixed Assets, Net Payables Accrued Expenses Long-Term Loan Common Equity 512566 367540 10000 5000 29408 1000 30000 61256 50000 11000 10000 50000 71256Assignment Notes: Answers are correct if they are within 5% of the correct response. most recent submissions to have. Grade appeals will be based on your Excel sheet . * Submit as many times as you would like before the submission deadline. I will take the » Use excel to solve these problem. It reduces rounding problems and is an important skill * Submitted answers should have two decimals places (i.e, 0.01). For things that are percentages, submit them as 2 whole numbers and two decimal places-23.20%. The margin for error on the grade is based off a two-decimal place calculation » Definitions for the ratios are on the course site (towards the bottom). Be careful, ROA, for example has several different definitions in the popular press so stick to these definitions rather than something you find through Google Accruals are debts that are soon due but have not yet been paid out - things like the salary owed to employees between pay periods show up as accruals » Assignment: Prepare an income statement and a balance sheet for is company using the information provided Calculate: 1. Current ratio 2. Quick ratio 3. NWC-to-total-Assets (Working capital to assets) 4. Ratio of total debt and liabilities to total assets 5. Ratio of total debt and liabilities to shareholders equity 6. Interest coverage 7. Net profit margin 8. Sales to total assets (Asset turnover) 9. Return on assets 10. Equity multiplieir 11. Return on equity
Answer #1

1 ) Current Ratio =Current Assets / Current Liabilities.

Current Assets = Cash + Receivables + Inventories = 1000+30000+61256 = $ 92256.

Current Liabilities = Payables + Accrued Expenses = 11000+10000 = $ 21000.

Current Ratio = 92256 / 21000 = 4.39 times.

Current Ratio indicates the ability of the firm to meet its current liabilities through current assets, ideally Current Ratio has to be around 1.25 but again varies from industry to industry, the higher the current ratio, the better.

2 ) Quick Ratio = Quick Assets / Current Liabilities.

Quick Assets = Cash + Receivables = 1000+30000 = $ 31000.

Quick Ratio = 31000 / 21000 = 1.47 times.

3 ) NWC to total assets = Net Working Capital / Total Assets

Net working capital = Current Assets - Current Liabilities = 92256-21000 = $ 71256

Total Assets = Cash + Receivables + Inventories + Fixed Assets = 1000+30000+61256 +50000 =$ 142256.

NWC to total assets = 71256/142256 = 50.09%.

4 ) Total Debt to Total Assets = Total debt / total assets.

Total Debt = Payables + Accrued Expenses + Longterm debt = 11000+10000+50000 = $ 71000.

Total Debt to Total Assets = 71000 / 142256 = 49.91%..